PHOTODISSOCIATION SPECTROSCOPY OF THE Ca^{+}-ACETYLENE π-COMPLEX

M. R. FRANCE, \underline{S. H. PULLINS}, and M. A. DUNCAN, Department of Chemistry, University of Georgia, Athens, GA 30602.

Ca^{+}-C_{2}H_{2} is produced in a molecular beam and studied with mass-selected photodissociation spectroscopy. A spectrum with sharp vibrational and rotational structure is measured near the Ca^{+} (^{2}P-^{2}S) transition. Ca^{+}-C_{2}H_{2} has a π-complex structure with a Ca^{+}-C_{2}H_{2} bond distance of r′′_{0} = 2.80 Å, a vibrational frequency (Ca^{+}-C_{2}H_{2} stretch) of ΔG′′_{1/2} = 169 cm^{−1} and a dissociation energy of D′′_{0} = 18.6 kcal/mol. This is the first determination of these properties for an isolated metal cation π-complex.